OpenStudy (anonymous):

A California couple found 1400 buried gold coins with a face value of $28,000, but a true value of $10,000,000. What is the added value per coin? Does this mean that each coin is worth $28,000 or all 1400 coins are worth $28,000. Cannot start to work out the problem until this is clear to me. Thanks.

OpenStudy (anonymous):

Each coin has a face value of $28,000 / 1400 coins = $20 Each coin has a face value of $20

OpenStudy (anonymous):

No that is not my question. My question is does the wording of the the question mean that the total value of the coins is 28K or is each coin valued at 28K. Thanks

OpenStudy (anonymous):

The total value of the coins is 28K

OpenStudy (anonymous):

Thanks! So I will 10,000,000/28,000 then subtract from that answer 20 to get the value added per coin.

OpenStudy (anonymous):

Yeah. I have an answer if you want to see if we get the same thing

OpenStudy (anonymous):

Ok I keep loosing my connection

OpenStudy (anonymous):

back on line!357-20=$337 added value each

OpenStudy (anonymous):

I get $7122.86 (10,000,000 / 28,000) * 20 = 7142.86 7142.86 - 20 = 7122.86

OpenStudy (anonymous):

Not close. Why did you multiply by $20? The original question wants to know how much value is added to the face value of each coin? If each coin is worth 20 face and the true value is 100,000 is it not closer to 357 each 357828,000=9996000

OpenStudy (anonymous):

I meant 10,000,000 I will try your way but I do not understand.

OpenStudy (anonymous):

this is for 6th grade

OpenStudy (anonymous):

We're trying to find the new value of the coin. The new value will be the ratio of the new value to the old value multiplied by the original value of the coin.

OpenStudy (anonymous):

But the total value must equal the true value of 10,000,000 and when I do it your way I get over 200,000,000

OpenStudy (anonymous):

You know what...i'm doing it the hard way. Just take 10,000,000 and divide by 1400. You get 7142.86

OpenStudy (anonymous):

Total value of all coins is 10,000,000 Per coin that's 10,000,000 divided by 1400 = 7142.86 So the added value per coin is 7142.86 - 20 = 7122.86

OpenStudy (anonymous):

That's easier to follow i think

OpenStudy (anonymous):

No value added is like the tax in Europe It adds to the original value.. The FAce value is the original value 20 dollar gold pieces. 81400=28,000

OpenStudy (anonymous):

Ok I see what you are saying. I think I will submit it both ways. Thanks so much

OpenStudy (anonymous):

28000/1400=20 each coins face value is 20 True value is 10,000,000-28,000=9972000/1400=7122.857143